OpenStudy (anonymous):

Annual profit in thousands of dollars is given by the function, P(x) = -.1x2 + 50x - 300, where x is the number of items sold, x ≥ 0. 1. describe the meaning of the number -.1 in the formula 2. describe the meaning of the number -300 in the formula 3. find the profit for 5 different values of x; PLEASE use x = 0, 100, 200, 350, and 500

OpenStudy (anonymous):

this looks like a quadratic. do u mean "-0.1x squared"?

OpenStudy (anonymous):

-300 are the fixed costs that occur even when nothing is sold

OpenStudy (anonymous):

and the -.1 corresponds to a market saturation when to much is sold, if I'm thinking in the right direction

OpenStudy (anonymous):

and part 3 is simple enough. just substitute the values and evaluate.

OpenStudy (anonymous):

yes I mean .1x^2

OpenStudy (anonymous):

can you do one substitute and evaluation for me? just so I know i am on track.....

OpenStudy (anonymous):

yup. just use the values for x they've given u, and work it out. like say, at x = 100 P(x) = -0.1(100)^2 +250(100) - 300 P(x) = -1000 + 25000 - 300 P(x) = 23700

OpenStudy (anonymous):

wouldnt -.1 times 100 give you 10 then 10^2 gives you 100?

OpenStudy (anonymous):

nope. work out the parentheses first, then the power of 2, and then the multiple outside the parentheses.

OpenStudy (anonymous):

so 100^2 then times-.1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you soooo much you are a saint saintdavid91 :-)